# Equal contribution

### v3

@@ Abstract
Music-driven dance generation has garnered significant attention due to its wide range of industrial applications, particularly in the creation of group choreography. During the group dance generation process, however, most existing methods still face three primary issues: multi-dancer collisions, single-dancer foot sliding and abrupt swapping in the generation of long group dance. In this paper, we propose TCDiff++, a music-driven end-to-end framework designed to generate harmonious group dance. Specifically, to mitigate multi-dancer collisions, we utilize a dancer positioning embedding to better maintain the relative positioning among dancers. Additionally, we incorporate a distance-consistency loss to ensure that inter-dancer distances remain within plausible ranges. To address the issue of single-dancer foot sliding, we introduce a swap mode embedding to indicate dancer swapping patterns and design a Footwork Adaptor to refine raw motion, thereby minimizing foot sliding. For long group dance generation, we present a long group diffusion sampling strategy that reduces abrupt position shifts by injecting positional information into the noisy input. Furthermore, we integrate a Sequence Decoder layer to enhance the model’s ability to selectively process long sequences. Extensive experiments demonstrate that our TCDiff++ achieves state-of-the-art performance, particularly in long-duration scenarios, ensuring high-quality and coherent group dance generation. Project Page.
@@ 1 Introduction
Figure 1: Visualizations of three key issues in baseline models: multi-dancer collisions [7], single-dancer foot sliding [57] and long group dance generation [23] where the blue man and the purple man suddenly swapped positions. In contrast, our approach eliminates these issues, delivering superior visual aesthetics.
@@
To address the aforementioned issues, we upgrade the preliminary conference version [7] into a fully end-to-end model. This end-to-end design not only produces more coherent body movements and positions but also significantly enhances performance in long-duration scenarios. Specifically, to mitigate multi-dancer collisions, we newly develop a Dancer Positioning Embedding (DPE) that incorporates spatial information by encoding features based on the relative left-right positioning of dancers. The DPE enables the model to better preserve dancer positioning patterns and effectively prevent collisions. Additionally, while the distance-consistency loss was originally applied only within the DTN module, in this work, we apply it to the entire end-to-end model to enforce global consistency. The end-to-end application of the distance-consistency loss leads to more effective control of the spacing between generated positions. To further address dancer ambiguity, we implement a simple yet effective Fusion Projection module (FP) to enhance the distinction between dancers in high-dimensional space, effectively mitigating dancer ambiguity. For single-dancer foot sliding, we newly integrate swap mode information, which specifies the start and end positions of dancers, thereby reducing spatial uncertainty caused by swap actions during generation. This enriched spatial information helps mitigate irregular foot sliding. Moreover, we refine the Footwork Adaptor plugin in TCDiff to directly adjust the generated results, improving the consistency between footwork and motion. To facilitate long dance generation, we remove the DTN module, which tends to cause oscillations due to spatial uncertainties in long-duration generation. To further reduce abrupt position shifts caused by dancer swapping, we propose the long group diffusion sampling strategy. This method constrains the input noise and incorporates positional priors, improving position consistency and minimizing sudden swaps. Additionally, we introduce the Sequence Decoder to strengthen the model’s sequence selection capability, further enhancing its performance in long-duration generation scenarios. In summary, our main contributions are:
  • To prevent multi-dancer collisions, we introduce a dancer positioning embedding to maintain relative positions and adapt the distance-consistency loss for end-to-end models, ensuring reasonable spacing. Additionally, a simple yet effective FP module is used to efficiently reduce dancer ambiguity.
@@ 2.1 Single-dancer Generation
Group dance generation extends single-dancer generation, but directly applying single-dancer methods to group dance presents significant challenges [19]. Early motion retrieval methods [22, 9, 41, 28] often produce deformed actions. Recent approaches use large datasets [26, 18, 50, 31, 32, 14] and deep learning techniques, such as auto-regressive models [2, 56, 1, 33, 45] and generative models [20, 48, 35, 34, 66] , to generate motions. Recently, diffusion-based models [16, 46, 48, 43] have emerged, achieving state-of-the-art performance with high diversity and fidelity. Solo dance generation emphasizes realism, where artifacts like foot sliding are unacceptable [58]. To address this, existing methods impose physical constraints through loss functions and foot contact labels [62, 64, 48, 35].
Since music often lasts for several minutes, and some musicals extend for hours, improving performance in long-duration scenarios is crucial. To address long-duration generation, some work [61, 35] proposes a progressive generation approach, where dance segments rich in choreography information are generated first, followed by a second stage that completes the transitions between these segments. The resulting segments are then concatenated to form a full-length dance sequence. However, single-dancer models often face issues with dancer ambiguity, arising from the severe similarity in dancer representations, result in multi-dancer collisions. In this paper, we address this issue by introducing a dancer positioning embedding, which incorporates relative left-right positioning information to aid spatial pattern learning. Furthermore, we directly apply the distance-consistency loss to the generated results, constraining the spacing between positions within a reasonable range.
@@ 2.2 Multi-dancer Generation
Multi-dancer generation is an emerging field in its early stages. To our knowledge, few studies [59, 24, 23, 57, 25, 7] have focused on scenarios with more than two dancers. Among them, GDanceR [24] and GCD [23] do not employ any specific structures to address motion representation imbalance, leading to dancer ambiguity. CoDancers [57] splits group motions into single-dancer motions to achieve scalability by incrementally adding dancers to the formation. As a result, this division causes the models to acquire only local information, leading to suboptimal group visual effects. TCDiff [7] is our previous version, a two-stage framework that first predicts dancers’ coordinates and then generates corresponding movements based on obtained coordinates. This separation preserves global formation features and ensures distinct trajectory coordinates, thus preventing dancer collisions caused by ambiguity. However, this two-stage generation process decouples footwork and movement modeling, making it prone to disjointed actions and displacements. Therefore, designing a model architecture that can both mitigate dancer ambiguity and simultaneously capture complete global and local information is in urgent need. In this paper, we upgrade the previous TCDiff to an end-to-end structure, enabling it to simultaneously capture global and local information. Furthermore, to mitigate dancer ambiguity, we modify the components from the original model and integrate new modules into the framework. These additions apply additional spatial constraints for further improvement.
@@ 2.3 Diffusion Model
In recent years, Diffusion Models (DM) [17, 47] have demonstrated outstanding performance in the content generation domain [5, 53, 54, 30, 29, 69]. Compared to earlier generative models [21, 10], such as Variational Auto-Encoders (VAE) [21] and Generative Adversarial Networks (GAN) [10], DMs offer significant advantages. Specifically, VAEs suffer from limited fidelity in learning complex, high-dimensional data distributions, while GANs are prone to mode collapse due to the challenges in balancing the adversarial relationship between generator and discriminator [55]. In contrast, DMs leverage probabilistic inference chains to learn complex distribution and do not have adversarial relationships, enabling higher fidelity and more stable training. Similar to previous generative approaches, diffusion models are also capable of conditional generation. This ability is particularly important in motion generation, especially for human motion and dance synthesis [36, 6]. In these tasks, conditional inputs, such as text [51, 68, 67], audio [3, 49] serve as crucial guidance, directing the model to generate the desired outcomes. However, dancer swapping introduces uncertainty during the diffusion sampling stage, leading to contradictory generation results. This contradiction leads to inconsistencies, as dancers’ positions vary across different sampling instances, resulting in abrupt swapping. To address this, we propose the long group diffusion sampling strategy, which constrains input noise and incorporates positional priors to enhance position consistency and reduce sudden swaps.
@@ 3.1 Problem Definition
Given an input music sequence =$\mathcal{M}=\left\{\boldsymbol{m}_{i}\right\}_{i=1}^{L}$ with $i=\left\{1,\dots,L\right\}$ indicates the index of the music frames, group choreography is to generate a corresponding group dance movement sequence $\boldsymbol{x}=\left\{\boldsymbol{x}^{(i)}\right\}_{i=1}^{L}$. Here, $\boldsymbol{x}^{(i)}=\left\{\boldsymbol{x}^{(i),c}\right\}_{c=1}^{C}$, where $\boldsymbol{x}^{(i),c}$ is the generated pose of $c$-th dancer at $i$-frame. $\boldsymbol{x}^{c}=\left\{\boldsymbol{x}^{(i),c}\right\}_{i=1}^{L}$ represents the movement sequence of the $c$-th dancer. For simplicity, we use $\boldsymbol{x}_{t}$ to represent $\{\boldsymbol{x}_{t}^{(i),c}\}$ at $t$ step.
@@ 3.2 Motion and Music Features
The dancer’s motion is modeled using a 24-joint SMPL framework  [37], with each joint’s pose $\boldsymbol{d}\in\mathbb{R}^{24\times 6=144}$ expressed in 6D rotation [65]. This model also includes binary contact indicators $\boldsymbol{f}\in\mathbb{R}^{4}$ for the heels and toes, and a 3D root position $\boldsymbol{p}\in\mathbb{R}^{3}$, forming a complete motion descriptor $\boldsymbol{x}=[\boldsymbol{f},\boldsymbol{p},\boldsymbol{d}]\in\mathbb{R}^{144+4+3=151}$. This rotation-based approach enhances motion smoothness over traditional 3D keypoint methods [63, 60, 39, 45]. For audio analysis, we utilize the Librosa toolkit [40] to derive a 35-dimensional feature set $\mathcal{M}\in\mathbb{R}^{35}$, which includes an audio envelope, 20 MFCCs for spectral details, 12 chroma features for pitch, and binary indicators for beats and peaks, essential for aligning movements with audio rhythms.
@@ 3.4 SSM Backbone
State Space Models (SSMs), such as structured state space sequence models [12] and Mamba [11, 8] are effective for long-sequence modeling due to their selective state transitions, linear computational complexity, and global receptive fields [52]. They map an input $x\left(t\right)\in\mathbb{R}$ to the output $y\left(t\right)\in\mathbb{R}$ through a hidden state $h\left(t\right)\in\mathbb{R}^{N}$, described by a linear ordinary differential equation (ODE):
$h^{\prime}\left(t\right)=\mathbf{A}h\left(t\right)+\mathbf{B}x\left(t\right),\quad y\left(t\right)=\mathbf{C}h\left(t\right).$ (3)
where $N$ denotes the state size, $\mathbf{A}\in\mathbb{R}^{N\times N}$ represents evolution parameter, $\mathbf{B}\in\mathbb{R}^{N\times 1}$ and $\mathbf{C}\in\mathbb{R}^{1\times N}$ are projection parameters. For practical computation, the Mamba discretize the system using zero-order hold (ZOH), yielding discrete parameters $\overline{\mathbf{A}}=\exp\left(\Delta\mathbf{A}\right)$, $\overline{\mathbf{B}}=\left(\Delta\mathbf{A}\right)^{-1}\left(\exp\left(\Delta\mathbf{A}\right)-\mathbb{I}\right)\Delta\mathbf{B}$, where the input-dependent step size $\Delta$ controls temporal granularity: small $\Delta$ resolves fine details (e.g., rapid motions), while large $\Delta$ captures long-term transitions. The discretized SSM computes outputs via global convolution with a structured kernel $\overline{\mathbf{K}}$:
$\overline{\mathbf{K}}$ $=\left(\mathbf{C}\overline{\mathbf{B}},\mathbf{C}\overline{\mathbf{A}\mathbf{B}},\dots,\mathbf{C}\overline{\mathbf{A}}^{L-1}\overline{\mathbf{B}}\right),$ (4)
@@
where $L$ is the input sequence length. This architecture improves the model’s selectivity, thereby enhancing its ability to process long sequences.
@@ 4.1 Group Dance Decoder
The Group Dance Decoder (GDD) reconstructs the input noisy dance motion sequence $\boldsymbol{x}_{T}$ into a clean raw motion sequence $\hat{\boldsymbol{x}}^{r}$. GDD uses conditional generation to obtain $\hat{\boldsymbol{x}}^{r}$, incorporating not only the music condition $\mathcal{M}$, but also the diffusion time step $T$ and the dancer spatial condition swap mode $S$ to guide the generation of dancer motions. Namely, the generation of GDD, denoted as $D\left(\cdot\right)$, executes a reverse diffusion process by estimating $D\left(\boldsymbol{x}_{T},T,\mathcal{M},S\right)\approx\boldsymbol{\hat{x}}^{r}$. To be more specific, GDD first incorporates the Dance Positioning Embedding (DPE) into the input features to enhance spatial representation. The Fusion Projection then processes the representation to reduce dancer ambiguity. Next, the Sequence Decoder enhances feature extraction through sequence selection, producing the final raw motion $\boldsymbol{\hat{x}}^{r}$. We provide a detailed introduction to the modules of GDD in the following sections, starting with the Dance Positioning Embedding (DPE).
Dance Positioning Embedding (DPE). We add the DPE to the input features to incorporate spatial information of the dancers, as shown in Figure 2. We find that dancers often display well-defined spatial relationships, such as maintaining a fixed left-right positional order to ensure no occlusion. Maintaining this positional information also helps alleviate dancer ambiguity and reduces the collisions. Based on the above findings, we designed the DPE to integrate additional spatial information into the features, helping to mitigate collisions. To be more specific, let the initial ground-truth motion sequences of $C$ dancers be denoted as $\left\{\boldsymbol{x}^{C}\right\}_{c=1}^{C}$, where each dancer has a x-axis position coordinate $p_{c}^{x-axis}\in\mathbb{R}$. These sequences are reordered from left to right based on positional coordinates using a permutation operator $\sigma(\cdot)$ to ensure monotonicity.
$\sigma=\operatorname{argsort}(p_{1}^{x-axis},p_{2}^{x-axis},\ldots,p_{C}^{x-axis}),$ (5)
$where,p_{\sigma(1)}^{x-axis}\leq p_{\sigma(2)}^{x-axis}\leq\cdots\leq p_{\sigma(N)}^{x-axis}.$ (6)
This yields the sorted sequence
$\boldsymbol{x}^{sorted}=\left\{\boldsymbol{x}^{\sigma(c)}\right\}_{c=1}^{C}.$ (7)
Next, the diffusion forward process is applied to $\boldsymbol{x}^{sorted}$ to inject Gaussian noise, resulting in the noisy input $\boldsymbol{x}^{sorted}_{T}$. Finally, a dance position embedding $\boldsymbol{DPE}$ is incorporated into each noisy sequence.
$\boldsymbol{x}_{T}=\boldsymbol{DPE}+\boldsymbol{x}_{T}^{sorted}.$ (8)
Here, $\boldsymbol{DPE}\in\mathbb{R}^{C}$ is a set of positional embeddings, which is automatically broadcasted to $\boldsymbol{DPE}^{\prime}\in\mathbb{R}^{C\times L\times d}$ during addition. Then, we input $\boldsymbol{x}_{T}$ into the Fusion Projection for further processing.
@@
The swap mode $S$ is an embedding derived from a sequence of numbers, representing the left-to-right positional order of the dancers in the final frame. This positional information comes from the left-to-right order of the final positions in the sorted sequence $\boldsymbol{x}^{sorted}$. The indices of this order are mapped through a linear transformation to generate the final swap mode $S$. Swap mode adds positional information, helping reduce spatial uncertainty caused by the dancer swapping during generation.

### v4

@@ Abstract
Music-driven dance generation has garnered significant attention due to its wide range of industrial applications, particularly in the creation of group choreography. During the group dance generation process, however, most existing methods still face three primary issues: multi-dancer collisions, single-dancer foot sliding and abrupt swapping in the generation of long group dance. In this paper, we propose TCDiff++, a music-driven end-to-end framework designed to generate harmonious group dance. Specifically, to mitigate multi-dancer collisions, we utilize a dancer positioning embedding to encode temporal and identity information. Additionally, we incorporate a distance-consistency loss to ensure that inter-dancer distances remain within plausible ranges. To address the issue of single-dancer foot sliding, we introduce a swap mode embedding to indicate dancer swapping patterns and design a Footwork Adaptor to refine raw motion, thereby minimizing foot sliding. For long group dance generation, we present a long group diffusion sampling strategy that reduces abrupt position shifts by injecting positional information into the noisy input. Furthermore, we integrate a Sequence Decoder layer to enhance the model’s ability to selectively process long sequences. Extensive experiments demonstrate that our TCDiff++ achieves state-of-the-art performance, particularly in long-duration scenarios, ensuring high-quality and coherent group dance generation. Project Page.
@@ 1 Introduction
Figure 1: Visualizations of three key issues in baseline models: multi-dancer collisions [8], single-dancer foot sliding [61], and long group dance generation [25], where the blue man and the purple man suddenly swapped positions. In contrast, our approach eliminates these issues, delivering superior visual aesthetics.
@@
To address the aforementioned issues, we upgrade the preliminary conference version [8] into a fully end-to-end model. This end-to-end design not only produces more coherent body movements and positions but also significantly enhances performance in long-duration scenarios. To be more specific, to mitigate multi-dancer collisions, we design DPE, which enriches feature representations with temporal and dancer identity information to reduce ambiguity and collisions. Additionally, while the distance-consistency loss was originally applied only within the DTN module, in this work, we apply it to the entire end-to-end model to enforce global consistency. The end-to-end application of the distance-consistency loss leads to more effective control of the spacing between generated positions. To further address dancer ambiguity, we implement a simple yet effective Fusion Projection module (FP) to enhance the distinction between dancers in high-dimensional space, effectively mitigating dancer ambiguity. For single-dancer foot sliding, we newly integrate swap mode information, which specifies the start and end positions of dancers, thereby reducing spatial uncertainty caused by swap actions during generation. This enriched spatial information helps mitigate irregular foot sliding. Moreover, we refine the Footwork Adaptor plugin in TCDiff to directly adjust the generated results, improving the consistency between footwork and motion. To facilitate long dance generation, we remove the DTN module, which tends to cause oscillations due to spatial uncertainties in long-duration generation. To further reduce abrupt position shifts caused by dancer swapping, we propose the long group diffusion sampling strategy. This method constrains the input noise and incorporates positional priors, improving position consistency and minimizing sudden swaps. Additionally, we introduce the Sequence Decoder to strengthen the model’s sequence selection capability, further enhancing its performance in long-duration generation scenarios. In summary, our main contributions are:
  • To prevent multi-dancer collisions, we introduce a dancer positioning embedding to to encode temporal and dancer identity information and adapt the distance-consistency loss for end-to-end models, ensuring reasonable spacing. Additionally, a simple yet effective FP module is used to efficiently reduce dancer ambiguity.
@@ 2.1 Single-dancer Generation
Group dance generation extends single-dancer generation, but directly applying single-dancer methods to group dance presents significant challenges [21]. Early motion retrieval methods [24, 10, 43, 30] often produce deformed actions. Recent approaches use large datasets [28, 20, 52, 33, 34, 16] and deep learning techniques, such as auto-regressive models [2, 59, 1, 35, 47] and generative models [22, 50, 37, 36, 68] , to generate motions. Recently, diffusion-based models [18, 48, 50, 45, 15, 60] have emerged, achieving state-of-the-art performance with high diversity and fidelity. Solo dance generation emphasizes realism, where artifacts like foot sliding are unacceptable [62]. To address this, existing methods impose physical constraints through loss functions and foot contact labels [65, 66, 50, 37].
Since music often lasts for several minutes, and some musicals extend for hours, improving performance in long-duration scenarios is crucial. To address long-duration generation, some works [64, 37, 38] adopt a progressive generation strategy, first producing choreography-rich dance segments and then generating transitions to seamlessly concatenate them into full-length sequences. However, single-dancer models often suffer from dancer ambiguity due to highly similar representations across dancers, leading to collisions in group choreography. To mitigate this, we propose a dancer positioning embedding (DPE) that encodes temporal and identity cues to distinguish dancers. In addition, we introduce a distance-consistency loss to keep inter-dancer spacing within reasonable limits.
@@ 2.2 Multi-dancer Generation
Multi-dancer generation is an emerging field in its early stages. To our knowledge, few studies [63, 26, 25, 61, 27, 8] have focused on scenarios with more than two dancers. Among them, GDanceR [26] and GCD [25] do not employ any specific structures to address motion representation imbalance, leading to dancer ambiguity. CoDancers [61] splits group motions into single-dancer motions to achieve scalability by incrementally adding dancers to the formation. As a result, this division causes the models to acquire only local information, leading to suboptimal group visual effects. TCDiff [8] is our previous version, a two-stage framework that first predicts dancers’ coordinates and then generates corresponding movements based on obtained coordinates. This separation preserves global formation features and ensures distinct trajectory coordinates, thus preventing dancer collisions caused by ambiguity. However, the two-stage generation process separates footwork from movement modeling, often leading to disjointed actions and unnatural displacements. Therefore, there is an urgent need for a model architecture that can reduce dancer ambiguity while jointly capturing both global and local information. In this paper, we upgrade the previous TCDiff into an end-to-end framework capable of learning global and local features simultaneously. To further mitigate dancer ambiguity, we modify the original components and introduce new modules that impose additional spatial constraints.
@@ 2.3 Diffusion Model
In recent years, Diffusion Models (DM) [19, 49] have demonstrated outstanding performance in the content generation domain [5, 56, 57, 32, 31, 71]. Compared to earlier generative models like VAEs [23] and GANs [11], diffusion models (DMs) offer clear advantages. VAEs often struggle with capturing complex, high-dimensional distributions, while GANs face instability and mode collapse due to their adversarial setup [58]. In contrast, DMs use probabilistic inference without adversarial training, leading to higher fidelity and more stable learning. Like previous methods, DMs also support conditional generation [6, 60]. This ability is particularly important in motion generation, especially for human motion and dance synthesis [39, 7]. In these tasks, conditional inputs, such as text [54, 70, 69] and audio [3, 51], guide the model to generate desired motions. However, dancer swapping causes inconsistencies in overlapping segments during diffusion sampling, resulting in abrupt transitions. To mitigate this, we introduce a long group diffusion sampling strategy that constrains noise and leverages positional priors to ensure consistent, long-duration generation.
@@ 3.1 Problem Definition
Given an input music sequence $\mathcal{M}=\left\{\boldsymbol{m}_{i}\right\}_{i=1}^{L}$ with $i=\left\{1,\dots,L\right\}$ indicates the index of the music frames, group choreography is to generate a corresponding group dance movement sequence $\boldsymbol{x}=\left\{\boldsymbol{x}^{(i)}\right\}_{i=1}^{L}$. Here, $\boldsymbol{x}^{(i)}=\left\{\boldsymbol{x}^{(i),c}\right\}_{c=1}^{C}$, where $\boldsymbol{x}^{(i),c}$ is the generated pose of $c$-th dancer at $i$-frame. $\boldsymbol{x}^{c}=\left\{\boldsymbol{x}^{(i),c}\right\}_{i=1}^{L}$ represents the movement sequence of the $c$-th dancer. For simplicity, we use $\boldsymbol{x}_{t}$ to represent $\{\boldsymbol{x}_{t}^{(i),c}\}$ at each timestep $t$ of the diffusion process.
@@ 3.2 Motion and Music Features
The dancer’s motion is modeled using a 24-joint SMPL  [40], with each joint’s pose $\boldsymbol{d}\in\mathbb{R}^{24\times 6=144}$ expressed in 6D rotation [67]. This model also includes binary contact indicators $\boldsymbol{f}\in\mathbb{R}^{4}$ for the heels and toes, and a 3D root position $\boldsymbol{p}\in\mathbb{R}^{3}$, forming a complete motion descriptor $\boldsymbol{x}=[\boldsymbol{f},\boldsymbol{p},\boldsymbol{d}]\in\mathbb{R}^{144+4+3=151}$. For audio, we use the Librosa [42] to build a music encoder that decomposes each audio clip into harmonic and percussive components and extracts mel-spectrograms, MFCCs (and deltas), chroma, onset/beat indicators, and a tempogram. The resulting features are concatenated and cached as fixed conditioning inputs.
@@ 3.4 SSM Backbone
State Space Models (SSMs) [13, 12, 9], have shown strong performance in modeling long sequences. Their effectiveness comes from three key features: selective control over internal state transitions, efficient linear-time computation, and the ability to capture global context across the entire sequence [55]. At their core, SSMs transform an input signal $x(t)\in\mathbb{R}$ into an output $y(t)\in\mathbb{R}$ through a hidden state $h(t)\in\mathbb{R}^{N}$, governed by a linear ordinary differential equation (ODE):
$h^{\prime}(t)=\mathbf{A}h(t)+\mathbf{B}x(t),\quad y(t)=\mathbf{C}h(t),$ (3)
where $N$ is the hidden state size, $\mathbf{A}\in\mathbb{R}^{N\times N}$ controls the state evolution, and $\mathbf{B}\in\mathbb{R}^{N\times 1}$, $\mathbf{C}\in\mathbb{R}^{1\times N}$ are input and output projection matrices. To support discrete-time computation, Mamba [12] discretizes the system using zero-order hold (ZOH), resulting in:
$\overline{\mathbf{A}}$ $=\exp(\Delta\mathbf{A}),$ (4)
$\overline{\mathbf{B}}$ $=(\Delta\mathbf{A})^{-1}(\exp(\Delta\mathbf{A})-\mathbb{I})\Delta\mathbf{B},$ (5)
where $\Delta$ is the input-dependent step size that controls temporal granularity: small $\Delta$ captures fine-grained dynamics (e.g., rapid motions), while large $\Delta$ captures long-term patterns.
The discretized model computes outputs through a global convolution with a structured kernel $\overline{\mathbf{K}}$:
$\overline{\mathbf{K}}$ $=(\mathbf{C}\overline{\mathbf{B}},\mathbf{C}\overline{\mathbf{A}\mathbf{B}},\dots,\mathbf{C}\overline{\mathbf{A}}^{L-1}\overline{\mathbf{B}}),$ (6)
@@
where $L$ is the input sequence length. This architecture allows the model to efficiently incorporate information from the entire sequence, improving its ability to handle long-range dependencies.
@@ 4.1 Group Dance Decoder
The Group Dance Decoder (GDD) reconstructs the input noisy dance motion sequence $\boldsymbol{x}_{T}$ into a clean raw motion sequence $\hat{\boldsymbol{x}}^{r}$. GDD uses conditional generation to obtain $\hat{\boldsymbol{x}}^{r}$, incorporating not only the music condition $\mathcal{M}$, but also the diffusion time step $T$ and the dancer spatial condition swap mode $S$ to guide the generation of dancer motions. Namely, the generation of GDD, denoted as $D\left(\cdot\right)$, executes a reverse diffusion process by estimating $D\left(\boldsymbol{x}_{T},T,\mathcal{M},S\right)\approx\boldsymbol{\hat{x}}^{r}$. To be more specific, GDD first incorporates the Dancer Positioning Embedding (DPE) into the input features to enhance spatial representation. The Fusion Projection then processes the representation to reduce dancer ambiguity. Next, the Sequence Decoder enhances feature extraction through sequence selection, producing the final raw motion $\boldsymbol{\hat{x}}^{r}$. We provide a detailed introduction to the modules of GDD in the following sections, starting with the Dancer Positioning Embedding (DPE).
Dancer Positioning Embedding (DPE) encodes temporal and identity information to uniquely represent each dancer across time. We first rearrange the clean ground-truth motion sequences such that dancers are ordered from left to right according to their horizontal (x-axis) positions in the initial frame, ensuring a consistent spatial layout across samples. After sorting, we apply the diffusion forward process to the rearranged motion sequences, injecting Gaussian noise, and obtain the noisy data $\boldsymbol{x}_{T}$. To encode temporal and dancer identity information into the noisy data $\boldsymbol{x}_{T}$, we first compute the DPE using the sinusoidal positional encoding from Transformer [53]. The positional encoding for any index $u$ (which can be either a time index $i$ or a dancer index $c$) at embedding dimension $2k$ or $2k+1$ is computed as:
$\mathrm{PE}(u,2k)=\sin\left(\frac{u}{10000^{\frac{2k}{d}}}\right),$ (7)
$\mathrm{PE}(u,2k+1)=\cos\left(\frac{u}{10000^{\frac{2k}{d}}}\right),$ (8)
where $k=0,1,\ldots,\frac{d}{2}-1$, and $d$ is the embedding dimension. For each dancer $c$ at frame $i$, the DPE vector is defined as:
$\boldsymbol{e}_{c,i}=\mathrm{PE}(i)+\mathrm{PE}(c).$ (9)
We integrate the DPE into the noisy motion data by performing an element-wise addition for each pose vector:
$\boldsymbol{x}_{T}^{\prime\,(i),c}=\boldsymbol{x}_{T}^{(i),c}+\boldsymbol{e}_{c,i}.$ (10)
This yields the final encoded noisy sequence:
$\boldsymbol{x}^{\prime}_{t}=\left\{\boldsymbol{x}_{t}^{\prime\,(i),c}\;\middle|\;i=1,\dots,L;\;c=1,\dots,C\right\}.$ (11)
This ensures that even after noise injection, the model retains both dancer-specific and temporal information, enhancing the consistency and controllability of generated motion sequences.
@@
Swap Mode (SM) identifies whether dancers change their relative positions between the start and end of each segment. Taking a 4-dancer group performance as an example, as illustrated in Fig. 2, we assign IDs 0, 1, 2, and 3 to the four dancers based on their initial positions along the x- and y-axes, ordered from left to right and top to bottom. Consequently, the initial x- and y-axis orders are always $[0,1,2,3]$. At the final frame, the new x- and y-axis orders are recorded and concatenated into a single index sequence, such as $[0,3,2,1,1,0,2,3]$. As shown in the figure, this index sequence clearly captures the position swap between the two dancers (indicated by the red arrow), effectively encoding the swap information. This sequence is subsequently embedded into a high dimension swap mode vector $\boldsymbol{S}\in\mathbb{R}^{d}$, which is then integrated with other conditioning inputs, including the music and the diffusion timestep.